# PrefixKV: Adaptive Prefix KV Cache is What Vision Instruction-Following Models Need for Efficient Generation

### v3

@@ 4.1 Experimental Settings
We follow [39] to employ LLaVA-1.5-7B [36] and LLaVA-1.5-13B [36], and leverage the LLaVA-Description [39] and MM-Vet [68] instruction-following datasets for evaluation. LLaVA-Description is a curated subset of 1000 detailed description instructions from the LLaVA-1.5 training set [39]. MM-Vet encompasses a diverse set of tasks designed to comprehensively evaluate the model performance in both understanding and generation. Besides, we also re-conduct the instruction tuning for LLaVA-1.5 models, to exclude the LLaVA-Description for preventing data leakage during evaluation. We follow [39] to employ the perplexity (PPL) and the ROUGE score [34] metrics. Specifically, PPL quantifies the exponential value of the cross-entropy loss between the predicted next token and ground truth. A lower PPL indicates the better generation quality. The ROUGE score calculates the longest common subsequence between the generated output and the reference outputs, with F1 score used for evaluation. A higher ROUGE score indicates the better consistency with reference responses.
Table 1: Comparison with SOTA methods on LLaVA-Description with the PPL / ROUGE metrics under various compression ratio budgets. Note that a lower PPL is better, while a higher ROUGE is better. The results without the KV compression are 3.20 / 0.62 for LLaVA-1.5-7B and 2.73 / 0.63 for LLaVA-1.5-13B, respectively. It can be observed that our method consistently achieves superior performance.
@@
Following [39], we utilize the model without cache compression to generate the reference output for the ROUGE score evaluation. Besides, in practice, multiple generation texts of a model can be different when the temperature is not zero. Thus, for the compression budget of 100%, i.e., without compression, we measure the ROUGE score of two generation outputs and it is thus smaller than 1. For other budgets, we use the temperature of 0 for the reproducibility and the ROUGE score may thus exceed the uncompressed one. We simply leverage 10 random samples from the training set of model for the global prefix configuration estimation offline. During decoding, we follow [39] to fix the distance to the latest token for pruning and maintain the layer-wise cache size ratios to satisfy the target budget. We compare with the state-of-the-art Elastic Cache [39], Heavy-Hitter Oracle [75], and StreamingLLM [61], which are termed as Elastic, H2O, and Local for brevity, respectively.
Table 3: Global prefix configuration. Uncompressed result is 5.28.
Method 10% 20% 30% 40% 50% 60% 70% 80% 90%
Baseline 41.8 26.6 20.4 15.4 11.8 9.06 6.47 5.75 5.72
Pyramid. 20.8 10.4 7.50 5.75 5.63 5.50 5.41 5.28 5.28
PrefixKV 7.38 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
@@ 4.2 Main Results
As shown in Table 1 and Table 2, our method consistently achieves the state-of-the-art performance compared with others across various compression budgets and different model scales. For example, as shown in Table 1, under a compression budget of 50% on LLaVA-Description, our PrefixKV significantly outperforms H2O and Elastic cache by 9.49 and 2.90 in PPL for LLaVA-1.5-7B. It also surpasses the Local cache by a notable margin of 0.39 in ROUGE score. We also observe that our advantages over others can be further amplified as the compression budget decreases. Under a compression budget of 20%, our PrefixKV can still maintain a satisfactory PPL of 3.69, which is reduced by 101.3, 44.6, and 10.3 compared with Local, H2O, and Elastic, respectively. For large model of LLaVA-1.5-13B, our method also demonstrates the notable superiority over others. For example, as shown in Table 2, our PrefixKV outperforms Elastic and H2O cache by 1.53 and 5.82 PPL, respectively, under a compression budget of 40% on MM-Vet. Furthermore, our method shows the minimal performance decline compared with the original model without KV compression in various scenarios. For example, as shown in Table 2, for LLaVA-1.5-13B, our method can maintain the nearly identical PPL and ROUGE scores compared with the uncompressed model at any compression budget exceeding 30%. Under a compression budget of 20%, our PrefixKV only leads to the decline of 0.44 PPL. These results well demonstrate the superiority of ours.
@@ 4.3 Model Analyses
We present comprehensive analyses for our method. Following [39], experiments are conducted on MM-Vet based on LLaVA-1.5-7B with PPL for evaluation, by default, .
Global prefix configuration matters. We verify the effectiveness of our method in identifying the ideal global prefix configuration. We first introduce the baseline, which keeps the same retained KV cache size for all layers. As shown in Table 3, our PrefixKV consistently brings performance benefit under various compression budgets. For example, under the compression budget of 30%, it surpasses the baseline by significant margin of 14.7 PPL. Besides, as the compression budget gradually decreases from 90% to 10%, its superiority becomes increasingly evident, showing increasing performance improvements. We also compare ours with PyramidKV [73], which manually allocates larger cache size in shallow layers and smaller size in deep layers. Since it compresses the cache only after the prefilling, we integrate it in the baseline for fair comparisons. As shown in Table 3, our strategy achieves superior performance over PyramidKV in various scenarios, especially under the low compression budget. These results show that compared with the same cache size across layers and the manually designed scheme by PyramidKV, our method can identify better global prefix configuration, which retains the overall contextual information more effectively.
Inference efficiency. We evaluate the inference efficiency of the model with our KV compression method to verify its benefit for acceleration. We follow [39] to conduct the evaluation in two practical scenarios. Specifically, firstly, we construct the input with 1024 prompt tokens and generate 512 tokens. Secondly, we employ a shorter input with 624 prompt tokens and generate 256 tokens, which means the minimal prompt length with only the image tokens and system prompts. We measure the inference time on the NVIDIA A100 GPU and use the batch size which maximizes the available memory to simulate the realistic deployment scenarios for better efficiency and throughput. As shown in Table 4, our method shows the notable inference speedups under various compression budgets compared with original model with full KV cache. For example, under the compression budget of 20% with the batch size of 16 for LLaVA-1.5-7B, our method demonstrates 1.8$\times$ inference speedup in terms of throughput. Besides, it also reduces the memory usage, avoiding OOM and enabling efficient inference with large batch size of 48. We also note that in this scenario, our method maintains the superior performance with only 0.69 PPL decline compared with the original model. These results well demonstrate the benefit of our method in practical deployment for efficient LVLMs.
@@
Table 5: Comparisons between offline estimation and online binary searching for each sample. Uncompressed result is 5.28.
@@
Offline 7.38 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
Online 7.38 5.97 5.66 5.53 5.50 5.41 5.38 5.28 5.28
Table 6: Impact of sample numbers. Uncompressed result is 5.28.
@@
Effect of offline estimation. Given a compression budget, we estimate the optimal KV cache size for each layer by random samples offline, which avoids the overhead of online searching. To verify its effectiveness, we first examine the variation of retained KV cache size ratios at each layer across different samples. As shown in Figure 4, different samples exhibit the similar cache size ratios for each layer. This indicates the potential of using the samples to estimate the optimal cache size ratios offline. We further present the comparison results between offline estimation and online binary searching for each sample in Table 5. It can be observed that offline estimation can obtain the comparable performance with the online searching. Therefore, our method can be integrated into the models efficiently, without the extra inference overhead. We also inspect the impact of adopting different numbers of random samples. As shown in Table 6, the performance is stable and robust across different sample sizes. Besides, employing a single sample can achieve good performance, which shows the effectiveness of offline estimation.
Relation between KV cache size and importance distribution. We derive better KV cache size for each layer based on the importance distribution. To provide deeper insights for their relation, we visualize the KV cache size ratios and the gini coefficients of priority sequence for different samples in Figure 4. We observe that layers with higher gini coefficients, i.e., with more concentrated importance distributions typically have fewer KV cache sizes, and vice versa. This qualitatively validates the reliability of our method, as layers with concentrated importance distributions can retain most information with fewer KV vectors, while layers with more dispersed importance distributions require larger KV cache sizes. Besides, we note that the retained KV cache size exhibits a W-shaped trend across layers. This inspires that utilizing more cost-effective attention mechanism in shallow to mid layers, as well as in mid to deep layers, may lead to more efficient model architecture.
Table 7: Eviction or merging. Uncompressed result is 5.28.
@@
Eviction or merging. Our PrefixKV shows favorable performance by cache eviction, i.e., retaining the prefix vectors and pruning the unimportant ones. To mitigate the information loss of eviction, previous research explores merging the less important KV vector with its nearest important one in position [39]. Therefore, we inspect the performance variation under combining our PrefixKV with cache merging. In addition to merging based on positional distance, we also examine the way based on the feature similarity, which are explored in the token merging field [5, 58]. Specifically, we denote the $l$-th layer’s retained KV vector set as $\Omega_{l}^{r}$ and pruned set as $\Omega_{l}^{p}$, respectively. The cache merging operation is the same for key and value vectors in different heads, we thus proceed with the key vector for illustration and omit the head superscript. For each pruned key vector $\boldsymbol{k}_{l}^{m}$ where $m\in\Omega_{l}^{p}$, we obtain its matching metric $\boldsymbol{c}_{l}^{mn}$ to each retained one $\boldsymbol{k}_{l}^{n}$ where $n\in\Omega_{l}^{r}$. For each $\boldsymbol{k}_{l}^{n}$, we obtain the set of pruned vectors that match with it by $\boldsymbol{T}_{l}^{n}=\{m\in\Omega_{l}^{p}|n=\text{argmax}_{u}(\boldsymbol{c}_{l}^{mu})\}$. The vectors are then merged to update $\boldsymbol{k}_{l}^{n}$ by $\boldsymbol{k}_{l}^{n*}=\text{Average}(\{\boldsymbol{k}_{l}^{u}|u\in\boldsymbol{T}_{l}^{n}\cup\{n\}\})$. We experiment with matching metrics based on the positional distance and feature cosine similarity by $\boldsymbol{c}_{l}^{mn}=-|m-n|$ and $\boldsymbol{c}_{l}^{mn}=\frac{\boldsymbol{k}_{l}^{m}\cdot\boldsymbol{k}_{l}^{n}}{||\boldsymbol{k}_{l}^{m}|||\boldsymbol{k}_{l}^{n}||}$, which are denoted as “Position” and “Feature”, respectively. As shown in Table 7, integrating “Position” can lead to inferior performance over PrefixKV in certain scenarios. We note that our method enables the maximal preservation of important KV vectors, and merging based on the positional distance, however, could introduce the interference to important cache due to the feature discrepancies among the vectors [58]. Besides, we observe that with less feature dispersion, combining “Feature” can bring the marginal improvements. This indicates that our method can well retain the significant contextual information across layers and eliminate the need for cache merging to reduce information loss, demonstrating its efficacy.
Figure 5: Mean absolute error between features of output tokens with and without compression across layers for 100 random samples. PrefixKV shows notably less feature disturbance than others.
Analyses for the feature disturbance. To further verify the effectiveness of our method in preserving valuable contextual information across layers, we inspect the feature perturbation caused by KV cache compression for output tokens at each layer. Specifically, for each output token, we calculate the mean absolute error between its feature with and without compression. We employ the compression budget of 50% and visual the average error across output tokens for samples. As shown in Figure 5, our method consistently exhibits lower mean absolute error between features compared with others. This shows that our method can introduce less interference to the features of output tokens, demonstrating enhanced retention of contextual information and resulting in improved generation quality.
Table 8: Comparisons on Qwen-VL. Uncompressed result is 6.28.
@@
Generalizability on other LVLMs. Following [39], we conduct experiments on Qwen-VL [2] to verify the general effectiveness of our method. In Table 8, our method consistently exhibits superior performance over others across various compression budgets, highlighting its generalizability.
Table 9: Chat generation example. It shows that others fail to generate reasonable responses while ours can ensure the generation quality with better contextual information retention across layers.
User What do you think is going on in this snapshot?
Local The two girls, two girls, two girls, two girls, two girls,…
H2O The image shows two young girls are two young girls are two young girls are two girls…
Elastic The image shows a young girls are two young girls are playing with a dog is a small children are playing with a small children are two young girls are playing with a toy dog.
Ours The image features two young girls standing next to each other, both holding stuffed animals. One girl is holding a teddy bear, while the other girl has a stuffed dog. They appear to be enjoying their time together.
Visualization of chat generation. In Table 9, we provide the generation example in real-word scenario. We set the compression budget to 20% and observe that previous methods all fail to answer the question, generating repetitive or confusing responses. In contrast, our method delivers coherent and rational outputs, demonstrating the superiority. More examples can be referred to the supplementary.

### v4

@@ 4.1 Experimental Settings
We follow liu2024efficient to employ LLaVA-1.5-7B liu2024improved and LLaVA-1.5-13B liu2024improved , and leverage the LLaVA-Description liu2024efficient and MM-Vet yu2023mm instruction-following datasets for evaluation. LLaVA-Description is a curated subset of 1000 detailed description instructions from the LLaVA-1.5 training set liu2024efficient . MM-Vet encompasses a diverse set of tasks designed to comprehensively evaluate the model performance in both understanding and generation. Besides, we also re-conduct the instruction tuning for LLaVA-1.5 models, to exclude the LLaVA-Description for preventing data leakage during evaluation. We follow liu2024efficient to employ the perplexity (PPL) and the ROUGE score lin2004rouge metrics. Specifically, PPL quantifies the exponential value of the cross-entropy loss between the predicted next token and ground truth. A lower PPL indicates the better generation quality. The ROUGE score calculates the longest common subsequence between the generated output and the reference outputs, with F1 score used for evaluation. A higher ROUGE score indicates the better consistency with reference responses.
Table 1: Comparison with SOTA methods on LLaVA-Description with the PPL / ROUGE metrics under various compression ratio budgets. Note that a lower PPL is better, while a higher ROUGE is better. The results without the KV compression are 3.20 / 0.62 for LLaVA-1.5-7B and 2.73 / 0.63 for LLaVA-1.5-13B, respectively. It shows that our method consistently achieves superior performance.
@@
Following liu2024efficient , we utilize the model without cache compression to generate the reference output for the ROUGE score evaluation. Besides, in practice, multiple generation texts of a model can be different when the temperature is not zero. Thus, for the compression budget of 100%, i.e., without compression, we measure the ROUGE score of two generation outputs and it is thus smaller than 1. For other budgets, we use the temperature of 0 for the reproducibility and the ROUGE score may thus exceed the uncompressed one. We simply leverage 10 random samples from the training set of model for the global prefix configuration estimation offline. During decoding, we follow liu2024efficient to fix the distance to the latest token for pruning and maintain the layer-wise cache size ratios to satisfy the target budget. We compare with the state-of-the-art Elastic Cache liu2024efficient , Heavy-Hitter Oracle zhang2024h2o , StreamingLLM xiao2023efficient , and PyramidKV zhang2024pyramidkv , which are termed as Elastic, H2O, Local, and Pyramid for brevity, respectively.
@@ 4.2 Main Results
As shown in Tab.˜1 and Tab.˜2, our method consistently achieves the state-of-the-art performance compared with others across various compression budgets and different model scales. For example, as shown in Tab.˜1, under a compression budget of 50% on LLaVA-Description, our PrefixKV significantly outperforms H2O and Elastic cache by 9.49 and 2.90 in PPL for LLaVA-1.5-7B. It also surpasses the Local cache by a notable margin of 0.39 in ROUGE score. We also observe that our advantages over others can be further amplified as the compression budget decreases. Under a compression budget of 20%, our PrefixKV can still maintain a satisfactory PPL of 3.69, which is reduced by 8.71, 44.6, and 10.3 compared with Pyramid, H2O, and Elastic, respectively. For large model of LLaVA-1.5-13B, our method also demonstrates the notable superiority over others. For example, as shown in Tab.˜2, our PrefixKV outperforms Elastic and H2O cache by 1.53 and 5.82 PPL, respectively, under a compression budget of 40% on MM-Vet. Furthermore, our method shows the minimal performance decline compared with the original model without KV compression in various scenarios. For example, as shown in Tab.˜2, for LLaVA-1.5-13B, our method can maintain the nearly identical PPL and ROUGE scores compared with the uncompressed model at any compression budget exceeding 30%. Under a compression budget of 20%, our PrefixKV only leads to the decline of 0.44 PPL. These results well demonstrate the superiority of ours.
@@ 4.3 Model Analyses
We present comprehensive analyses for our method. Following liu2024efficient , experiments are conducted on MM-Vet based on LLaVA-1.5-7B with PPL for evaluation, by default, .
Inference efficiency.
[…]
Global prefix configuration matters.
We verify the effectiveness of our method in identifying the ideal global prefix configuration. We first introduce the baseline, which keeps the same retained KV cache size for all layers. As shown in Tab.˜4, our PrefixKV consistently brings performance benefit under various compression budgets. For example, under the compression budget of 30%, it surpasses the baseline by significant margin of 14.7 PPL. Besides, as the compression budget gradually decreases from 90% to 10%, its superiority becomes increasingly evident, showing increasing performance improvements. We also compare ours with PyramidKV zhang2024pyramidkv , which manually allocates larger cache size in shallow layers and smaller size in deep layers. Since it compresses the cache only after the prefilling, we integrate it in the baseline for fair comparisons. As shown in Tab.˜4, our strategy achieves superior performance over PyramidKV in various scenarios, especially under the low compression budget. These results show that compared with the same cache size across layers and the manually designed scheme by PyramidKV, our method can identify better global prefix configuration, which retains the overall contextual information more effectively.
@@
Table 4: Prefix config. (Uncompressed: 5.28).
@@
Baseline 41.8 26.6 20.4 15.4 11.8 9.06 6.47 5.75 5.72
Pyramid 20.8 10.4 7.50 5.75 5.63 5.50 5.41 5.28 5.28
PrefixKV 7.38 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
Table 5: Sample numbers. (Uncompressed: 5.28).
@@
Given a compression budget, we estimate the optimal KV cache size for each layer by random samples offline, which avoids the overhead of online searching. To verify its effectiveness, we first examine the variation of retained KV cache size ratios at each layer across different samples. As shown in Fig.˜4, different samples exhibit the similar cache size ratios for each layer. This indicates the potential of using the samples to estimate the optimal cache size ratios offline. We further present the comparison results between offline estimation and online binary searching for each sample in Tab.˜6. It can be observed that offline estimation can obtain the comparable performance with the online searching. Therefore, our method can be integrated into the models efficiently, without the extra inference overhead. We also inspect the impact of adopting different numbers of random samples. As shown in Tab.˜5, the performance is stable and robust across different sample sizes. Besides, employing a single sample can achieve good performance, which shows the effectiveness of offline estimation.
Robustness of offline estimation.
To inspect the robustness of offline estimation, we measure the standard deviation (std) of the retained KV cache size ratios obtained via online binary search across all samples, as well as the mean absolute difference (mad) between these values and the configuration of offline estimation. We present the results in Tab.˜7. It quantitatively shows that the variance of retained KV cache size ratios across samples is small, indicating the feasibility of offline estimation. The small mean absolute difference supports the efficacy of offline estimation. We also conduct experiments using random samples from different domains and with different instruction complexities to further verify the robustness. Specifically, for different domains, we leverage random samples from open-knowledge visual question answering (VQA), optical character recognition (OCR), and complex reasoning (reasoning), respectively. In Tab.˜8, we present the PPL results. We observe that the performance remains stable across different domains. Besides, for instruction complexities, we obtain random samples with different instruction lengths: shorter than 100, between 100 and 500, and longer than 500. We present the results in Tab.˜9. It shows that our method also exhibits robustness to the instruction complexity, showing favorable generalizability. For the binary search for offline estimation, we also present analyses for its sensitivity and convergence behavior. Specifically, we conduct experiments under the compression ratio of 50% using varying approximation threshold $\delta$ to inspect the robustness of binary search. We present the average search steps and PPL results in Tab.˜10. It shows that the search consistently converges within 7-12 steps, with stable performance across different settings. We also fix the number of search steps of binary search for evaluation. As shown in Tab.˜11, the performance remains stable across difference scenarios, showing the reliability.
Table 6: Offline estimation.(Uncompressed: 5.28).
Method 10% 20% 30% 40% 50% 60% 70% 80% 90%
Offline 7.38 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
Online 7.38 5.97 5.66 5.53 5.50 5.41 5.38 5.28 5.28
Table 7: Deviation analyses.
10% 20% 30% 40% 50% 60% 70% 80% 90%
std 0.010 0.021 0.023 0.024 0.023 0.020 0.013 0.011 0.006
mad 0.010 0.021 0.024 0.024 0.023 0.020 0.014 0.011 0.006
Table 8: Robustness to domain.
Domain 10% 20% 30% 40% 50% 60% 70% 80% 90%
VQA 7.45 5.97 5.72 5.53 5.50 5.41 5.38 5.28 5.28
OCR 7.38 6.01 5.72 5.53 5.53 5.41 5.38 5.28 5.28
reasoning 7.40 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
Table 9: Robustness to instruction complexity.
Length 10% 20% 30% 40% 50% 60% 70% 80% 90%
(0, 100] 7.38 6.01 5.66 5.53 5.53 5.44 5.38 5.28 5.28
(100, 500] 7.42 5.97 5.72 5.53 5.50 5.44 5.38 5.28 5.28
(500, $\inf$) 7.38 5.97 5.72 5.54 5.50 5.41 5.38 5.28 5.28
Table 10: Robustness to $\delta$
$\delta$ 0.0125 0.025 0.05 0.075 0.1
avg. step 11.5 10.8 9.3 8.5 7.8
PPL 5.50 5.50 5.50 5.51 5.53
Table 11: Robust. to step
steps 7 8 9 10 11 12
PPL 5.53 5.51 5.51 5.50 5.50 5.50
Table 12: Merging. (Uncompressed: 5.28).
@@
Relation between KV cache size and importance distribution.
We derive better KV cache size for each layer based on the importance distribution. To provide deeper insights for their relation, we visualize the KV cache size ratios and the gini coefficients of priority sequence for different samples in Fig.˜4. We observe that layers with higher gini coefficients, i.e., with more concentrated importance distributions typically have fewer KV cache sizes, and vice versa. This qualitatively validates the reliability of our method, as layers with concentrated importance distributions can retain most information with fewer KV vectors, while layers with more dispersed importance distributions require larger KV cache sizes. Besides, we note that the retained KV cache size exhibits a W-shaped trend across layers. This inspires that utilizing more cost-effective attention mechanism in shallow to mid layers, as well as in mid to deep layers, may lead to more efficient model architecture. Besides, we also observe that the distributions of retained KV cache size ratios and Gini coefficients obtained from different samples exhibit low variance, indicating favorable consistency in importance patterns. The empirical stability further verifies the reliability of offline estimation using few samples, as they can approximate the overall layer-wise importance distribution.
Eviction or merging.
@@
Our PrefixKV shows favorable performance by cache eviction, i.e., retaining the prefix vectors and pruning the unimportant ones. To mitigate the information loss of eviction, previous research explores merging the less important KV vector with its nearest important one in position liu2024efficient . Therefore, we inspect the performance variation under combining our PrefixKV with cache merging. In addition to merging based on positional distance, we also examine the way based on the feature similarity, which are explored in the token merging field bolya2022token ; wei2023joint . Specifically, we denote the $l$-th layer’s retained KV vector set as $\Omega_{l}^{r}$ and pruned set as $\Omega_{l}^{p}$, respectively. The cache merging operation is the same for key and value vectors in different heads, we thus proceed with the key vector for illustration and omit the head superscript. For each pruned key vector $\boldsymbol{k}_{l}^{m}$ where $m\in\Omega_{l}^{p}$, we obtain its matching metric $\boldsymbol{c}_{l}^{mn}$ to each retained one $\boldsymbol{k}_{l}^{n}$ where $n\in\Omega_{l}^{r}$. For each $\boldsymbol{k}_{l}^{n}$, we obtain the set of pruned vectors that match with it by $\boldsymbol{T}_{l}^{n}=\{m\in\Omega_{l}^{p}|n=\text{argmax}_{u}(\boldsymbol{c}_{l}^{mu})\}$. The vectors are then merged to update $\boldsymbol{k}_{l}^{n}$ by $\boldsymbol{k}_{l}^{n*}=\text{Average}(\{\boldsymbol{k}_{l}^{u}|u\in\boldsymbol{T}_{l}^{n}\cup\{n\}\})$. We experiment with matching metrics based on the positional distance and feature cosine similarity by $\boldsymbol{c}_{l}^{mn}=-|m-n|$ and $\boldsymbol{c}_{l}^{mn}=\frac{\boldsymbol{k}_{l}^{m}\cdot\boldsymbol{k}_{l}^{n}}{||\boldsymbol{k}_{l}^{m}|||\boldsymbol{k}_{l}^{n}||}$, which are denoted as “Position” and “Feature”, respectively. As shown in Tab.˜12, integrating “Position” can lead to inferior performance over PrefixKV in certain scenarios. We note that our method enables the maximal preservation of important KV vectors, and merging based on the positional distance, however, could introduce the interference to important cache due to the feature discrepancies among the vectors wei2023joint . Besides, we observe that with less feature dispersion, combining “Feature” can bring the marginal improvements. This indicates that our method can well retain the significant contextual information across layers and eliminate the need for cache merging to reduce information loss, demonstrating its efficacy.
Figure 5: Mean absolute error vis.
@@
Table 13: Results on Qwen-VL. (Uncompressed: 6.28).
Analyses for the feature disturbance.
To further verify the effectiveness of our method in preserving valuable contextual information across layers, we inspect the feature perturbation caused by KV cache compression for output tokens at each layer. Specifically, for each output token, we calculate the mean absolute error between its feature with and without compression. We employ the compression budget of 50% and visual the average error across output tokens for 100 random samples across layers. As shown in Fig.˜5, our method consistently exhibits lower mean absolute error between features compared with others. This shows that our method can introduce less interference to the features of output tokens, demonstrating enhanced retention of contextual information and resulting in improved generation quality.
Generalizability on other LVLMs.
Following liu2024efficient , we conduct experiments on Qwen-VL bai2023qwen to verify the general effectiveness of our method. In Tab.˜13, our method consistently exhibits superiority over others across various compression budgets, highlighting generalizability.
Table 14: Results on LLMs.
Dataset Base. Local H2O Pyramid Ours
LongB. 41.6 32.0 35.0 37.1 37.5
Needle. 100 30.4 49.3 97.3 97.6
Table 15: With quant.
Method 12.5% 25% 50% 100%
KIVI 21.8 41.2 41.6 41.6
Ours 40.2 41.1 41.5 41.6
Comparison results on LLMs.
Our method can also be adopted for the KV cache compression in LLMs. Following zhang2024pyramidkv , we conduct experiments on LongBench bai2023longbench and Needle-in-a-HayStack liu2024lost ; fu2024data based on LLaMA-3-8B dubey2024llama . We set the KV cache size budget to 128 and present the results in Tab.˜14. We can see that our PrefixKV also outperforms others in such scenarios. It well reserves the valuable information for model’s generation in limited KV cache budget but under long context. The results further exhibit the general efficacy of our method. Besides, we also compare ours with the KV cache quantization methods for LLMs. We present the comparison results with advanced KIVI liu2024kivi on LongBench for LLaMA-3-8B in Tab.˜15. We find that in long context scenarios, for different budgets, KV cache favors different compression dimensions, i.e., precision in quantization and length in pruning. For extreme 12.5% compression budget, the length dimension shows more redundancy and our method performs better. For larger budgets, removing precision redundancy preserves more information and KIVI obtains slightly better performance. Moreover, our pruning method is complementary to quantization method and can be combined to eliminate redundancy in both dimensions. For example, combining KIVI under 4 bit with ours can achieve competitive performance of 40.0 with 3.125% compression ratio budget, leading to stronger efficacy.